An aqueous solution of 6.3 g of oxalic acid dehydrate is made up to 250 ML. The volume of  0.1 N  NAOH  required  to completely naturalise 10 ML of this solution is :

Grade:12th pass

1 Answers

Tony
5 years ago

Equivalent mass of C2H2O4.2H2O = 126/2 = 63

Normality of oxalic acid = 6.3/63 * 1000/250 = 0.4 N

We know that,

N1V1 = N2V2

Where N1 is normality of oxalic acid

Vis volume of oxalic acid

N2 is normality of sodium hydroxide

V2 is volume of sodium hydroxide

Therefore,

0.4 * 10 = 0.1 * V2

V2 = 40 ml.

Therefore 40 ml of NaOH is required to completely neutralize 10 ml of oxalic acid.